OpenStudy (kirbykirby):

Does the series converge or diverge? \[\sum_{1}^{infinity}\frac{\sqrt{n}}{n+1}\]

OpenStudy (jamesj):

Notice that \[ \frac{\sqrt{n}}{n+1} > \frac{1}{n+1} \] Now does the infinite sum of the terms on the right converge?

OpenStudy (kirbykirby):

Oh ok the right one diverges (similar to the harmonic one).. so the left one diverges by comparison test right?

OpenStudy (jamesj):

yes

OpenStudy (kirbykirby):

ok thanks :)!

OpenStudy (anonymous):

jamesj is exactly right, but with a little practice you do this with your eyeballs. degree of numerator is 1/2 and denominator is 1 if the degree of the denominator does not exceed the degree of the numerator by more than one it diverges. i will be willing to bet that is what jamesj thought (correct me if i am wrong) before he wrote the answer

OpenStudy (kirbykirby):

our prof won't allow us to "eyeball" stuff.. we have to show exactly each step -_-

OpenStudy (jamesj):

sat73 is right; I know that heuristic and used it here first. It's a convenient rule to have. Knowing now what the answer 'should be', I decided what was the best way to prove it. The method I suggested and you're now using is my estimation definitely the most elementary.